OpenStudy (anonymous):

Pls help with this... A and B are roots of x^2-(k+6)x+(2)(2k-1). Find value of k if A+B=(1/2)(AB)

OpenStudy (anonymous):

if there are two roots then\[(x-A)(x-B)=x ^{2}-(A+B)x+AB\]That being said,\[k+6=A+B\]\[2(2k-1)=AB\]Considering that\[A+B=(AB)/2\]\[2(2k-1)=AB \rightarrow 2k-1= AB/2\]We conclude that\[2k-1=k+6\rightarrow k= 7/3\] Cheers!

OpenStudy (anonymous):

@Syderitic Thanks!! I started same way but made some silly mistake in calculation!!! So, thanks for your help.......Medal for you....☺

OpenStudy (anonymous):

shouldn't A+B = -(k+6) ???

OpenStudy (anonymous):

@Syderitic I think u made a mistake in the last step... 2k-1 = k + 6 2k - k = 6 + 1 k = 7

OpenStudy (anonymous):

bummer you are right, sorry, I'll try and be more carefull next time. Problably hould hit the sack already, it's 6am in here. Oh well...

OpenStudy (anonymous):

Yeah, tired mind does play tricks!!!!☺